OpenStudy (anonymous):

The longer leg of a 30°-60°-90° triangle is 18. What is the length of the other leg

OpenStudy (tkhunny):

You need to know these. The legs are related by a factor of \(\sqrt{3}\). Brush up on this triangle and the 45-45-90 triangle. VERY IMPORTANT!

OpenStudy (anonymous):

The hypotenuse of a 30-60-90 triangle is twice as long as the shorter leg, so 2 • 18 = 36

OpenStudy (anonymous):

the answer choices are \[9\sqrt{3}\],9\[6\sqrt{3}\], and \[12\] by the way

OpenStudy (tkhunny):

Which of these is related to a longer leg of 18 by a factor of \(\sqrt{3}\)?

OpenStudy (tkhunny):

Dear Future, please reread the Code of Conduct, particularly premise #2 http://openstudy.com/code-of-conduct